OpenStudy (hopeless):

How to rewrite without rational exponents y 1/7

OpenStudy (anonymous):

can you draw the problem, its kind of confnusing

OpenStudy (hopeless):

y1/7

OpenStudy (anonymous):

\[(\sqrt[7]{y})^1=\sqrt[7]{y}\]

OpenStudy (hopeless):

that is all i have it says to rewrite without rational exponents

OpenStudy (anonymous):

\[\Large {y^{1/7}} = \sqrt[7]{{{y^1}}}\]

OpenStudy (anonymous):

we can use this symbol to indicate that something i rasied to something else: ^ exp. x^2, is x squared

OpenStudy (anonymous):

The basic rule for exponents is something to the effect of: \[\Large a \cdot {x^{n/m}} = a \cdot \sqrt[m]{{{x^n}}}\]